OpenStudy (anonymous):

find the inverse of f(x) =12 - (1/2)x

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

f^-1: x=12-(1/2)y solve for y x-12=-(1/2)y And now im stuck... LOL someone help xD

OpenStudy (anonymous):

is this \[f(x)=12-\frac{1}{2}x\] or \[f(x)=12-\frac{1}{2x}\]

OpenStudy (anonymous):

to calculate inverse , swap the x and f(x), the find f(x) in terms of x. u get the inverse. my answer is 2(12-x)

OpenStudy (anonymous):

first one \[x=12-\frac{1}{2}y\] \[x-12=-\frac{1}{2}y\] \[2x-24=-y\] \[y=24-2x\]

OpenStudy (anonymous):

thank you all!